Chad has a total of 5 people in his family.
Sally is Chad's twin sister
Chad's sister Hailey is 4 years older than Chad
Chad's mother was 24 when Chad was born
Chad's father is 3 years older than his mother
1. Write an expression for each family member's age as a function of Chad's age (x). Clearly label each expression. Show or explain all of your work even if you use mental math or a calculator.
2. The average age of the 5 family members is 22. Determine the age of Chad's sister Hailey. Show or explain all of your work even if you use mental math or a calculator.

Answer by haileytucki(390)

Chad = x
Sally = x
Hailey = x + 4
Mum = 24 + x
Dad = 27 + x
Total family age = 5x + 55
Average = total family age / Number of family members
22 = (5x + 55)/5
22 * 5 = 5x + 55
110 = 5x + 55
110 - 55 = 5x
55 = 5x
x = 11
Chad is 11, Sally is 11, Hailey is 15, Mum is 35, Dad is 38